OpenStudy (anonymous):

A solution is made by dissolving 25.5 grams of glucose (C6H12O6) in 398 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m?

OpenStudy (ookawaiioo):

You ready?

OpenStudy (ookawaiioo):

I had to write this one out....., which I rarely do

OpenStudy (ookawaiioo):

You gonna learn today

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Okay

OpenStudy (ookawaiioo):

Molality (m) = grams of solutes in/ 1000g of water

OpenStudy (ookawaiioo):

25.5 grams of glucose (C6H12O6) in 398 grams of water 25.5*1000/398 = m

OpenStudy (ookawaiioo):

Tf = -(tc)* m tc = -1.86 *m tc = freezing-point depression

OpenStudy (anonymous):

m = 64.0703 tc = -119.045

OpenStudy (ookawaiioo):

uhhhh let me calculate it

OpenStudy (ookawaiioo):

oops i skipped a step!

OpenStudy (ookawaiioo):

25.5 grams of glucose (C6H12O6) in 398 grams of water 25.5*1000/398 = x x/1000 = m

OpenStudy (anonymous):

m = 0.0640 tc = -0.119 tf = 0.00761

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@oOKawaiiOo

OpenStudy (ookawaiioo):

Winner winner chicken dinner

OpenStudy (anonymous):

So is that the answer then?

OpenStudy (ookawaiioo):

Your answer is tf

OpenStudy (anonymous):

okay, thanks!